Could I get some guidance on this please?
Consider the hypothesis test given by
H_0: μ=650
H_a: μ<650
Assume the population is normally distributed. In a random sample of 25 subjects, the sample mean is found to be x ̅=640, and the sample standard deviation is s=30.
(a) Is this test for population proportion, mean or standard deviation? What distribution should you apply for the critical value?
(b) Is the test a right-tailed, left-tailed or two-tailed test?

(c) Find the test statistic. (Show work and round the answer to two decimal places)







(d) Determine the P-value for this test. (Show work and round the answer to three decimal places if necessary. If you use technology to find the P-value, you have to describe the steps)





(e) Is there sufficient evidence to justify the rejection of at the α=0.05 level? Explain.
Thank you for your help!

Answer by ewatrrr(24785)


Hi
n = 25, x̄ = 640, s = 30
Ho: u = 650
Ha: u < 650
a) mean
We are using the T distribution because
1) we do not know the population standard deviation (very common in most cases)
2) n < 30
(b) Is the test a right-tailed, left-tailed or two-tailed test? left-tailed
(c) t = (640-650)/(30/5) = -10/6 = -5/3 = -1.6667
(d)P(t < -1.6667), DF 24 = .0543
(d) .0543 > .05
there is not sufficient evidence to justify the rejection of Ho at the α=0.05 level